convert 20.399 to the fractional equivalent and reduc
Answer #1
let
x = 20.399
10x = 203.99
100x = 2039.99
100x - 10x = 2039.99 - 203.99
90x = 1836
x = 1836/90 if 9 is coming again and again after 3 after decimal place
Answer #2
There is not much that can be done to figure out how to write 20.399 as a fraction, except to literally use what the decimal portion of your number, the .399, means.
Since there are 3 digits in 399, the very last digit is the "1000th" decimal place.

So we can just say that .399 is the same as 399/1000.



So your final answer is: 20.399 can be written as the fraction (20)399/1000
Answer #3
20399/1000